# Enhanced Single-Photon Detector: Achieving Superconducting-Level Performance with Conventional Quantum Technology

### v1

@@ 3 Discussion
1 (4,2) $(80.2\%,\ 1.2\times 10^{-3})$ (6,3) $(74.2\%,\ 4.5\times 10^{-5})$ (7,3) $(81.6\%,\ 7.2\times 10^{-5})$
2 (4,2) $(94.5\%,\ 5.9\times 10^{-5})$ (6,3) $(89.1\%,\ 7.3\times 10^{-8})$ (7,3) $(95.2\%,\ 1.8\times 10^{-7})$
3 (4,2) $(97.9\%,\ 2.3\times 10^{-5})$ (6,3) $(95.7\%,\ 1.1\times 10^{-7})$ (7,3) $(97.0\%,\ 2.4\times 10^{-7})$
4 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(96.8\%,\ 1.4\times 10^{-7})$ (7,3) $(97.2\%,\ 2.5\times 10^{-7})$
5 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(96.9\%,\ 1.4\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
6 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
7 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
8 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (7,3) $(97.2\%,\ 2.6\times 10^{-7})$
Para 4 Para 5 Para 6
[…]
1 (8,4) $(69.6\%,\ 1.8\times 10^{-6})$ (8,2) $(95.2\%,\ 4.2\times 10^{-3})$ (8,2) $(95.2\%,\ 4.2\times 10^{-3})$
2 (8,4) $(82.9\%,2.6\times 10^{-10})$ (8,4) $(95.6\%,\ 1.4\times 10^{-7})$ (8,3) $(97.2\%,\ 1.6\times 10^{-5})$
3 (8,4) $(92.5\%,5.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.8\%,\ 1.0\times 10^{-9})$
4 (8,4) $(95.2\%,8.2\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.7\%,9.4\times 10^{-10})$
5 (8,4) $(95.6\%,9.1\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
6 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
7 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
8 (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$ (8,4) $(95.6\%,9.3\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 3: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.81$.
@@
Table 4: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$, and high controlled gate fidelity, $P=0.99$.
@@
1 (4,1) $(78.6\%,\ 2.2\times 10^{-3})$ (6,1) $(87.6\%,\ 3.3\times 10^{-3})$ (8,1) $(92.3\%,\ 4.4\times 10^{-3})$
2 (4,2) $(97.5\%,\ 1.2\times 10^{-4})$ (6,2) $(98.7\%,\ 4.7\times 10^{-4})$ (8,4) $(96.7\%,\ 1.6\times 10^{-7})$
3 (4,2) $(98.9\%,\ 2.7\times 10^{-5})$ (6,3) $(98.0\%,\ 3.7\times 10^{-7})$ (8,4) $(96.9\%,\ 9.7\times 10^{-10})$
4 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
5 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4)) $(96.9\%,\ 9.8\times 10^{-10})$
6 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
7 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
8 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) $(97.9\%,\ 1.5\times 10^{-7})$ (8,4) $(96.9\%,\ 9.8\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 4: Performance scaling for initial parameters $\eta_{0}=27.5\%$, $d_{0}=1.0\times 10^{-6}$, with $P=0.99$.
Table 5: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$, and moderate controlled gate fidelity, $P=0.81$.
Para 1 Para 2 Para 3
[…]
1 (4,1) $(72.4\%,\ 2.2\times 10^{-3})$ (6,1) $(82.3\%,\ 3.3\times 10^{-3})$ (8,1) $(88.1\%,\ 4.4\times 10^{-3})$
2 (4,2) $(90.7\%,\ 1.1\times 10^{-4})$ (6,2) $(97.7\%,\ 4.6\times 10^{-4})$ (8,4) $(94.4\%,\ 1.5\times 10^{-7})$
3 (4,2) $(97.3\%,\ 2.3\times 10^{-5})$ (6,3) $(97.0\%,\ 3.2\times 10^{-7})$ (8,4) $(95.5\%,\ 8.8\times 10^{-10})$
4 (4,2) $(98.2\%,\ 2.3\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
5 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4)) $(95.6\%,\ 9.3\times 10^{-10})$
6 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
7 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
8 (4,2) $(98.3\%,\ 2.4\times 10^{-5})$ (6,3) $(97.0\%,\ 1.5\times 10^{-7})$ (8,4) $(95.6\%,\ 9.3\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 5: Performance scaling for initial parameters $\eta_{0}=27.5\%$, $d_{0}=1.0\times 10^{-6}$, with $P=0.81$.
Table 6: Performance evolution of an ESPD initialized with a baseline detector from [28], with $p$ varied from 0.86 to 0.98. Here, $P=0.99$ and experiments are run for $(n,k)=(5,2)$ on each level.
pp=0.86 pp=0.89 pp=0.92 pp=0.95 pp=0.98
1 $(70.5\%,1.8\times 10^{-3})$ $(75.3\%,1.8\times 10^{-3})$ $(80.4\%,1.8\times 10^{-3})$ $(85.8\%,1.8\times 10^{-3})$ $(91.4\%,1.8\times 10^{-3})$
2 $(76.3\%,1.3\times 10^{-4})$ $(82.6\%,1.3\times 10^{-4})$ $(88.4\%,1.4\times 10^{-4})$ $(93.4\%,1.5\times 10^{-4})$ $(97.6\%,1.6\times 10^{-4})$
3 $(78.6\%,2.8\times 10^{-5})$ $(84.8\%,3.3\times 10^{-5})$ $(90.0\%,3.8\times 10^{-5})$ $(94.3\%,4.2\times 10^{-5})$ $(97.9\%,4.6\times 10^{-5})$
4 $(79.5\%,2.6\times 10^{-5})$ $(85.4\%,3.0\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
5 $(79.9\%,2.6\times 10^{-5})$ $(85.6\%,3.0\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
6 $(80.0\%,2.6\times 10^{-5})$ $(85.6\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
7 $(80.0\%,2.7\times 10^{-5})$ $(85.6\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
8 $(80.0\%,2.7\times 10^{-5})$ $(85.7\%,3.1\times 10^{-5})$ $(90.3\%,3.4\times 10^{-5})$ $(93.4\%,3.7\times 10^{-5})$ $(97.9\%,4.0\times 10^{-5})$
(a) Detection efficiency $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
(b) Dark count rate $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
Figure 6: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.99$.
@@
1 (6,2) $(64.0\%,\ 2.4\times 10^{-3})$ (7,2) $(69.8\%,\ 3.3\times 10^{-3})$ (8,2) $(74.5\%,\ 4.2\times 10^{-3})$
2 (6,2) $(68.0\%,\ 2.6\times 10^{-4})$ (7,2) $(78.5\%,\ 5.5\times 10^{-4})$ (8,2) $(85.5\%,\ 1.1\times 10^{-3})$
3 (6,2) $(71.5\%,\ 4.2\times 10^{-5})$ (7,2) (84.4%, $4\%,\ 1.0\times 10^{-4}) (8$,2) $(90.6\%,\ 2.3\times 10^{-4})$
4 (6,2) $(74.7\%,\ 3.3\times 10^{-5})$ (7,2) (87.7%, $7\%,\ 6.8\times 10^{-5}) (8$,2) $(92.3\%,\ 1.2\times 10^{-4})$
5 (6,2) $(77.4\%,\ 3.5\times 10^{-5})$ (7,2) $(89.4\%,\ 7.0\times 10^{-5})$ (8,2) $(92.9\%,\ 1.1\times 10^{-4})$
6 (6,2) $(79.5\%,\ 3.8\times 10^{-5})$ (7,2) $(90.1\%,\ 7.3\times 10^{-5})$ (8,2) $(93.0\%,\ 1.1\times 10^{-4})$
7 (6,2) $(81.0\%,\ 4.0\times 10^{-5})$ (7,2) $(90.4\%,\ 7.4\times 10^{-5})$ (8,2) $(93.1\%,\ 1.1\times 10^{-4})$
8 (6,2) $(82.1\%,\ 4.2\times 10^{-5})$ (7,2) $(90.5\%,\ 7.5\times 10^{-5})$ (8,2) $(93.1\%,\ 1.1\times 10^{-4})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 7: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
3.1 Device Requirements
The flexibility of the ESPD paradigm, which allows for various DOF to be employed, enables the selection of an easily manipulated parameter. In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates. Specifically, $Q$ can be suppressed to below the $10^{-4}$ level using commercial platforms[41, 42], thereby guaranteeing condition $Q\ll 1$.
The primary technological requirements of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$. In practice, the transmission rate $p$ is often approximated as $p\approx 1$, representing a high transmission efficiency [43, 44, 45, 46, 47, 48, 49, 42]. However, since $p$ decays exponentially as the number of stages $n$ increases, experimental characterization of $p$ remains crucial. Table 6 and Fig.6 demonstrate the performance when $p$ is degraded, showing that the ESPD framework remains valid even if $p$ drops to $0.86$, albeit with reduced performance compared to higher values.
Attention now shifts to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities, exceeding $99\%$ [39, 40] using modern technology and reaching $81\%$ via linear-optical methods [38], suggesting that high values of $P$ are readily attainable. Our numerical results in Section 2.4 confirm that the ESPD paradigm does not rely on an excessively demanding value for $P$, showing similar substantial performance gains for $P=0.99$ and $P=0.81$. Furthermore, Table 7 and Fig.7 illustrate the robustness of the scheme, confirming its efficacy even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This demonstrated robustness confirms that the technological demands of the ESPD method are fully compatible with existing quantum technologies.
It is also important to note that the implementation of the ESPD scheme does not require perfect auxiliary photon sources. The parameter $P$ relies solely on the detection of the state $|1\rangle$ from the auxiliary path when the control signal input is non-vacuum. This condition holds true regardless of the exact number of photons in the auxiliary path or the presence of unwanted $|0\rangle$ measurement outcomes, which are inherently filtered out [^2] This justifies the necessity of the dedicated filter block within the design.. Intuitively, multi-photon sources might even offer practical benefits over single-photon sources by providing more opportunities for detecting the required final $|1\rangle$, although a rigorous theoretical formulation of this potential advantage is reserved for future work.

### v2

@@ 2.4 Performance Simulations
1 (4,1) $(76.9\%,\ 2.2\times 10^{-3})$ (6,1) (85.5%, $5\%,\ 3.4\times 10^{-3}) (8$,1) $(90.0\%,\ 4.5\times 10^{-3})$
2 (4,2) $(95.3\%,\ 1.2\times 10^{-4})$ (6,2) $(94.8\%,\ 3.8\times 10^{-4})$ (8,4) $(93.1\%,\ 1.7\times 10^{-7})$
3 (4,2) $(97.7\%,\ 2.6\times 10^{-5})$ (6,3) $(95.5\%,\ 1.4\times 10^{-7})$ (8,4) $(93.3\%,\ 8.4\times 10^{-10})$
4 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.4\times 10^{-10})$
5 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
6 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
7 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
2.5 Application on QKD
To illustrate the practical impact of the ESPD scheme, we consider its application in QKD. For a QKD protocol having a secure threshold $e_{th}$ and total error rate $e_{C}$ not accounting for SPDs, the minimal tolerable channel transmission rate $\gamma$ is conventionally approximated by:
Given that the ESPD paradigm reduces the DCR ($d$) by several orders of magnitude (even compared to SNSPDs), it proportionately lowers the minimal tolerable channel transmission rate $\gamma$ by the same factor. Furthermore, the simultaneous enhancement of the DE ($\eta$) not only lowers $\gamma$ but also translates directly into an improved QKD key rate at any fixed distance.
Compared to the empty-signal-detection paradigm [27], which also reduces the threshold of the minimal tolerable channel transmission rate, the ESPD scheme avoids introducing an extra sifting term that penalizes efficiency. This allows significant QKD performance improvements without additional efficiency costs. Consequently, the ESPD paradigm offers a compelling first-choice solution for practical quantum communication implementations.
3 Discussion
We now discuss the technological feasibility, as well as limitations and avenues for future work.
3.1 Technological Feasibility and Performance under Degraded Parameters
The most significant advantage of the ESPD paradigm lies in its reliance on accessible, room-temperature components, effectively bypassing the cryogenic requirements of superconducting SPDs. The device requirements are discussed as follows.
3.1.1 State Preparations and Measurements
The flexibility of the architecture, which allows for various DOF to be employed, enables the selection of an easily manipulated one. In practical implementations, this facilitates state preparation and projective measurement, ensuring that the error parameter $Q$ can be executed with extremely low rates. Specifically, $Q$ can be suppressed to $10^{-4}$ level using standard optical setups and commercial platforms [43, 38], thereby guaranteeing $Q\ll 1$.
3.1.2 Controlled Gates
The primary technological prerequisites of the ESPD scheme are the controlled quantum gates, typically C-NOT operations, which define the critical parameters $p$ and $P$.
A high transmission efficiency, with attenuation as low as $0.1$ dB (corresponding to a propagation probability $p\approx 0.98$) [38], is achievable with commercial photonic integrated circuits [44, 45, 46, 47], suggesting that $p$ can approach unity. It is important to reiterate that $p$ represents the intrinsic transmission rate of a single, non-cascaded controlled module, and consequently, does not decrease with an increasing number of cascaded modules or cascaded levels. The cumulative system loss is fully accounted for in the derivation of the total system efficiency, $\eta_{s+1}$ (Eq.(6)).
However, to accommodate practical implementations, one may consider using post-selection controlled gates instead of the feedforward configurations. While the ESPD design remains valid for such probabilistic operations, the overall system efficiency is necessarily degraded. Specifically, if the post-selection efficiency of the controlled gate is $N$, the average transmission rate from one effective module to the next will be $p^{\frac{1}{N}}$. Although a full theoretical calculation for the post-selection case is structurally similar to that established in Section 2.2 and is not repeated here, the resulting degradation underscores the importance of characterizing system performance under reduced transmission rate. Table 4 and Fig.4 demonstrate that the ESPD framework maintains performance enhancement even when $p$ drops to $0.80$ (approximately 1 dB attenuation, corresponding to the loss cascaded about 10 controlled modules with 0.1 dB attenuation on each.). This result indicates robustness, validating the scheme’s feasibility even for C-NOT gates with a low post-selection efficiency of around $10\%$, which is well within the reach of current photonic technology [48, 49, 50, 51, 52, 38].
(a) DE $\eta_{s}$ vs. $p$ for $(n_{s},k_{s})=(5,2)$.
(b) DCR $d_{s}$ vs. level $p$ for $(n_{s},k_{s})=(5,2)$.
Figure 4: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$. All other settings remain consistent with those described in Section 2.4.
Table 4: Performance evolution of an ESPD initialized with a baseline detector from [28], with $p$ varied from 0.80 to 0.96. Experiments are run for $(n,k)=(5,2)$ on each level, and all other settings remain consistent with those described in Section 2.4.
pp=0.80 pp=0.84 pp=0.88 pp=0.92 pp=0.96
1 $(60.8\%,1.8\times 10^{-3})$ $(66.7\%,1.8\times 10^{-3})$ $(72.9\%,1.8\times 10^{-3})$ $(79.7\%,1.8\times 10^{-3})$ $(87.0\%,1.8\times 10^{-3})$
2 $(61.0\%,1.1\times 10^{-4})$ $(70.6\%,1.2\times 10^{-4})$ $(79.7\%,1.3\times 10^{-4})$ $(87.8\%,1.4\times 10^{-4})$ $(94.6\%,1.5\times 10^{-4})$
3 $(61.0\%,1.8\times 10^{-5})$ $(72.6\%,2.4\times 10^{-5})$ $(82.1\%,3.1\times 10^{-5})$ $(89.5\%,3.7\times 10^{-5})$ $(95.4\%,4.3\times 10^{-5})$
4 $(60.9\%,1.5\times 10^{-5})$ $(73.5\%,2.2\times 10^{-5})$ $(82.8\%,2.8\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
5 $(60.8\%,1.5\times 10^{-5})$ $(74.0\%,2.2\times 10^{-5})$ $(83.1\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
6 $(60.8\%,1.5\times 10^{-5})$ $(74.2\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
7 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
8 $(60.8\%,1.5\times 10^{-5})$ $(74.3\%,2.3\times 10^{-5})$ $(83.2\%,2.9\times 10^{-5})$ $(89.9\%,3.4\times 10^{-5})$ $(95.4\%,3.8\times 10^{-5})$
We now turn our attention to the parameter $P$, which is intrinsically linked to the fidelity of the C-NOT gate. Notably, $P$ is lower-bounded by the gate’s overall fidelity, as $P$ quantifies deviation only on the controlled path, whereas the full C-NOT fidelity accounts for the correct operation rate of both the control and controlled paths. Contemporary C-NOT gates demonstrate high fidelities without low-temperature requirement, suggesting that high values of $P$ are readily attainable[48, 49, 50, 51, 52, 38]. Specifically, the recent feedforward photonic C-NOT experiment can achieve fidelity over 97% [39, 40, 41, 42]. However, it is also worth testing the effectiveness of the design under degraded parameters to verify its suitability for standard fabrication processes with lower technology requirements. Here, we also provide numerical results for two degraded scenarios: $P=0.80$ (see Table 5 with visual representation in Fig.5, and $P=0.40$, (see Table 6 with visual representation in Fig.6). All other settings remain consistent with those described in Section 2.4. The numerical results confirm the robustness of the scheme: Comparison of Table 5 to Table 2 demonstrates that the ESPD paradigm does not rely on an excessively demanding value for $P$, since similar substantial performance gains are shown for both $P=0.97$ and $P=0.80$, while Table 6 implies that the scheme is effective even when the C-NOT fidelity parameter is severely degraded to $P=0.40$. This demonstrated robustness indicates that the technological demands of the ESPD method are fully compatible with existing non-superconducting quantum technologies.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 5: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.80$.
@@ 3.1.2 Controlled Gates
Table 5: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.80$.
@@
1 (4,2) $(77.8\%,\ 1.2\times 10^{-3})$ (6,2) $(88.5\%,\ 2.4\times 10^{-3})$ (8,2) $(92.3\%,\ 4.2\times 10^{-3})$
2 (4,2) $(91.3\%,\ 5.7\times 10^{-5})$ (6,3) $(92.4\%,\ 2.1\times 10^{-6})$ (8,4) $(91.0\%,\ 1.4\times 10^{-7})$
3 (4,2) $(95.7\%,\ 2.2\times 10^{-5})$ (6,3) $(93.3\%,\ 1.3\times 10^{-7})$ (8,4) $(90.6\%,\ 7.6\times 10^{-10})$
4 (4,2) $(96.4\%,\ 2.3\times 10^{-5})$ (6,3) $(93.5\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
5 (4,2) $(96.5\%,\ 5.5\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
6 (4,2) $(96.5\%,\ 2.4\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
7 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
8 (4,2) $(96.5\%,\ 2.3\times 10^{-5})$ (6,3) $(93.6\%,\ 1.3\times 10^{-7})$ (8,4) $(90.5\%,\ 7.5\times 10^{-10})$
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for different $(n_{s},k_{s})$.
Figure 6: Performance for parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.40$.
@@
1 (2,1) $(75.1\%,\ 3.2\times 10^{-2})$ (7,2) $(66.8\%,\ 3.3\times 10^{-3})$ (8,2) $(71.1\%,\ 4.2\times 10^{-3})$
2 (6,2) $(78.2\%,\ 2.1\times 10^{-2})$ (7,2) $(72.8\%,\ 5.4\times 10^{-4})$ (8,2) $(79.6\%,\ 1.0\times 10^{-3})$
3 (6,2) $(79.1\%,\ 9.5\times 10^{-3})$ (7,2) $(77.2\%,\ 9.0\times 10^{-4})$ (8,2) $(84.3\%,\ 2.1\times 10^{-4})$
4 (6,2) $(78.7\%,\ 2.4\times 10^{-3})$ (7,2) $(80.1\%,\ 5.7\times 10^{-5})$ (8,2) $(86.4\%,\ 1.0\times 10^{-4})$
5 (6,2) (77.7%, $7\%,\ 2.9\times 10^{-5}) (7$,2) $(82.0\%,\ 5.8\times 10^{-5})$ (8,2) $(87.2\%,\ 9.5\times 10^{-5})$
6 (6,2) $(76.7\%,\ 5.4\times 10^{-5})$ (7,2) $(83.0\%,\ 6.1\times 10^{-5})$ (8,2) $(87.5\%,\ 9.5\times 10^{-5})$
7 (6,2) $(76.0\%,\ 3.8\times 10^{-5})$ (7,2) $(83.6\%,\ 6.3\times 10^{-5})$ (8,2) $(87.6\%,\ 9.6\times 10^{-5})$
8 (6,2) $(75.4\%,\ 3.7\times 10^{-5})$ (7,2) $(83.9\%,\ 6.4\times 10^{-5})$ (8,2) $(87.7\%,\ 9.6\times 10^{-5})$
3.1.3 Choice of initialized SPDs